Find the minimum and maximum values of f(x,y,z)=8x^2-2y subject to x^2+y^2=1


1
Expert's answer
2022-03-01T15:32:50-0500

We are asked for the extreme values of "f" subject to the constraint "g(x,y,z)=x^2+y^2=1."

Using Lagrange multipliers, we solve the equations "\nabla f(x,y)=\lambda\nabla g(x, y)" and "g(x,y)=1,"

which can be written as


"f_x=\lambda g_x, f_y=\lambda g_y, x^2+y^2=1"

"16x=\lambda(2x), -2=\lambda(2y), x^2+y^2=1"

We have "x=0" or "\lambda=8."

If "x=0," then "y=-1" or "y=1."

If "\lambda=8," then "y=-1\/8" and "x=\pm\dfrac{3\sqrt{7}}{8}"

"f(0, -1)=8(0)^2-2(-1)=2"

"f(0, 1)=8(0)^2-2(1)=-2"

"f(-\dfrac{3\sqrt{7}}{8}, -\dfrac{1}{8})=8(\dfrac{-3\sqrt{7}}{8})^2-2(-\dfrac{1}{8})=\dfrac{65}{8}"

"f(\dfrac{3\sqrt{7}}{8}, -\dfrac{1}{8})=8(\dfrac{3\sqrt{7}}{8})^2-2(-\dfrac{1}{8})=\dfrac{65}{8}"

Therefore the maximum value of on the circle "x^2+y^2=1" is "f(\pm\dfrac{3\sqrt{7}}{8}, -\dfrac{1}{8})=\dfrac{65}{8} ," and the

minimum value is "f(0, 1)=-2."